OpenStudy (anonymous):

Solve this quadratic y^2-2y-4y = 0 what are the roots? Please show all working

OpenStudy (anonymous):

y^2 -6y = 0 y(y-6) = 0 y = 0, y = 6

OpenStudy (anonymous):

\[y^2-6y=0\][ \[y(y-6)=0\] \[y=0\] \[y-6=0\] \[y=6\] \[\{0,6\}\]